# DynoSurf: Neural Deformation-based Temporally Consistent Dynamic Surface Reconstruction

### v1

@@ 4.3 Ablation Study
5 Conclusion and Discussion
Figure 11: Failure case of our method.
Although our method has achieved good performance on many examples, it may encounter limitations in cases where the point cloud in certain areas is sparse or the local structure is elongated but exhibits significant non-rigid deformations (as depicted in the Fig. 11). This is attributed to our utilization of Chamfer distance as the alignment supervision, which restricts the establishment of correspondences between adjacent frames solely based on spatial proximity. Incorporating more robust metrics for shape alignment would enhance the effectiveness of our approach.
Supplementary Material
More visual results
We present more visual results in this section and showcase the complete action sequences in the supplementary video.
Comparisons with state-of-the-art methods.  We present additional visual results for comparisons with LPDC [50] and Cadex [21] on the DFAUST dataset [3] (Fig. 12), AMA dataset [52] (Fig. 13) and DT4D dataset [25] (Fig. 15). It is evident that our method outperforms other approaches.
@@ 5 Conclusion and Discussion
Refer to caption(a) Input(b) LPDC [50](c) Cadex [21](d) Ours00.15
Refer to caption(a)(b)(c)(d)(e)(f)
Figure 14: Comparison of visual results by different variants of excluding a certain loss in our method. (a) Input; (b) w.o. $L_{\textrm{R-SDF}}$; (c) w.o. $L_{\textrm{Norm}}$; (d) w.o. $L_{\textrm{Smo}}$; (e) w.o. $L_{\textrm{Shape}}$; (f) Ours.
Refer to caption(a) Input(b) Cadex [21](c) Ours00.0600.200.1
Ablation study.  Additionally, Fig. 14 illustrates the visual results of our method when different loss terms are omitted. We can see that our method, incorporating all loss terms, achieves the best results.
Performance of our method on noisy data and partially missing data.  Furthermore, in Fig. 16, we showcase more visual results of our method applied to noisy data constructed from the DFAUST dataset [3] and partially missing data constructed from AMA dataset [52], respectively.
@@
Technical Details
In this section, we explain the definitions for some loss functions and evaluation metrics mentioned in the paper. The Chamfer Distance between $\mathbf{X}$ and $\mathbf{Y}$ is defined as
$\texttt{CD}_{\ell_{p}}(\mathbf{X},\mathbf{Y})=\frac{1}{|\widetilde{\mathbf{X}}|}\sum_{\mathbf{x}\in\widetilde{\mathbf{X}}}\left\|\mathbf{x}-\hat{\mathbf{y}}\right\|_{\ell_{p}}+\frac{1}{|\widetilde{\mathbf{Y}}|}\sum_{\mathbf{y}\in\widetilde{\mathbf{Y}}}\left\|\mathbf{y}-\hat{\mathbf{x}}\right\|_{\ell_{p}},$ (15)
where $p=1,2$. $\hat{\mathbf{y}}\in\widetilde{\mathbf{Y}}$ and $\hat{\mathbf{x}}\in\widetilde{\mathbf{X}}$ are the closest points for the $\mathbf{x}$ and $\mathbf{y}$ respectively. When $\mathbf{X}(\mathbf{Y})$ is a point set, $\widetilde{\mathbf{X}}=\mathbf{X}(\widetilde{\mathbf{Y}}=\mathbf{Y})$. When $\mathbf{X}(\mathbf{Y})$ represents a mesh, $\widetilde{\mathbf{X}}(\widetilde{\mathbf{Y}})$ denotes the sampling point on the mesh. In the learning phase, we set the number of sampling points to $10^{4}$, and during the evaluation, we set the number of sampling points to $10^{5}$.
In Eqs. (3) and (10), the normal consistency of $\mathbf{X}$ and ${\mathbf{Y}}$ is defined as
$\texttt{NC}_{\ell_{1}}(\mathbf{X},\mathbf{Y})=\frac{1}{|\mathbf{N}_{x}|}\sum_{\mathbf{n}_{x}\in\mathbf{N}_{x}}\left|1-|\langle\mathbf{n}_{x},\mathbf{n}_{\hat{y}}\rangle|\right|+\frac{1}{|\mathbf{N}_{y}|}\sum_{\mathbf{n}_{y}\in\mathbf{N}_{y}}\left|1-|\langle\mathbf{n}_{y},{\mathbf{n}}_{\hat{x}}\rangle|\right|,$ (16)
@@
$\textrm{Recall}(\mathbb{M}_{k},\mathbb{M}_{k}^{\textrm{GT}},\epsilon)=\left|\left\{\mathbf{v}_{1}\in\widetilde{\mathbf{V}}_{k},\textrm{s.t.}\min_{\mathbf{v}_{2}\in\widetilde{\mathbf{V}}_{k}^{\textrm{GT}}}\|\mathbf{v}_{1}-\mathbf{v}_{2}\|<\epsilon\right\}\right|,$
$\textrm{Precision}(\mathbb{M}_{k},\mathbb{M}_{k}^{\textrm{GT}},\epsilon)=\left|\left\{\mathbf{v}_{2}\in\widetilde{\mathbf{V}}_{k}^{\textrm{GT}},\textrm{s.t.}\min_{\mathbf{v}_{1}\in\widetilde{\mathbf{V}}_{k}}\|\mathbf{v}_{1}-\mathbf{v}_{2}\|<\epsilon\right\}\right|.$
@@
We denote $\{\mathbf{V}^{\textrm{GT}}(k)\}_{k}$ as the temporally consistent corresponding points on the ground-truth surfaces. We set $|\mathbf{V}^{\textrm{GT}}(k)|=10^{5}$. In the first frame, we construct the index set of nearest points:
$\texttt{Index}=\{\rho_{i}|\mathbf{v}_{\rho_{i}}=\mathop{\arg\min}_{\mathbf{v}\in\mathbf{V}_{1}}\|\mathbf{v}-\mathbf{v}_{i}^{\textrm{GT}}(1)\|\},$
where $\mathbf{V}_{1}$ is the vertex set of $1$-st reconstructed mesh $\mathbb{M}_{1}$, $\mathbf{v}_{i}^{\textrm{GT}}(1)\in\mathbf{V}^{\textrm{GT}}(1)$. Following [34, 21], the Correspondences Error is defined as

### v2

@@ Appendix 0.A Supplement of ablation study
0.A.0.2 Arbitrarily long point cloud sequence.
Our method is theoretically independent of the sequence length. We experimented with sequences of various lengths on "50007__\__shake__\__shoulders" sequence from DFAUST. As shown in Tab. 3 below, our method can effectively handle sequences of varying lengths. When the sequence is relatively long, the reconstruction quality decreases slightly due to the diversity of deformations.
Table 3: Results of our method on sequences with various lengths.
Sequence length CD(×$\times 10^{-4}$) $\downarrow$ NC $\uparrow$ F-$0.5\%$uparrow F-$1\%$uparrow Corr. (×$\times 10^{-2}$) $\downarrow$
10 0.154 0.957 0.866 0.992 0.53
30 0.138 0.954 0.876 0.994 0.97
50 0.167 0.955 0.831 0.989 1.46
80 0.265 0.946 0.704 0.960 1.72
0.A.0.3 Various data qualities.
We experimented with point clouds of different qualities on the AMA dataset: (1) we randomly sampled $N_{k}$ points from the ground truth mesh respectively as input of our method; and (2) we randomly sampled $5000$ points from ground truth mesh, randomly selected $10$ points from them, and deleted the $h_{k}$ nearest neighbor points of these points to construct point clouds with holes. As listed in Tab. 4, our method shows robustness to variations in point cloud density and holes. We also showcase the visualization of constructed point clouds with various qualities in Fig. 11.
Table 4: Results of our method on point clouds with various qualities.
Input Data CD(×$\times 10^{-4}$) $\downarrow$ NC $\uparrow$ F-$0.5\%$uparrow F-$1\%$uparrow Corr. (×$\times 10^{-2}$) $\downarrow$
(1) $N_{k}=2000$ 0.451 0.906 0.563 0.899 3.54
$N_{k}=3000$ 0.402 0.911 0.610 0.916 3.46
$N_{k}=4000$ 0.381 0.913 0.597 0.921 3.41
(2) $h_{k}=10$ 0.336 0.918 0.634 0.935 2.88
$h_{k}=20$ 0.366 0.914 0.617 0.926 3.57
$h_{k}=30$ 0.405 0.911 0.587 0.912 3.50
GT-normal 0.318 0.918 0.656 0.940 3.10
Ours 0.324 0.919 0.649 0.939 2.89
[…]
Figure 11: Visualization of an input point cloud with various qualities.
0.A.0.4 Dependence on the normal direction.
In all experiments, we did not use any ground-truth normals. Instead, we first use the algorithm in PyMeshLab [33] to estimate the normals of input point clouds. We then adjust the normal directions to ensure that the majority of normals point outward. Besides, we also replaced estimated normals with ground-truth normals in our method. As listed in Tab. 4, the performance is comparable to Ours, demonstrating the robustness of our method to normals of input point clouds.
0.A.0.5 Visualization of omitting loss terms.
Additionally, Fig. 12 illustrates the visual results of our method when different loss terms are omitted. We can see that our method, incorporating all loss terms, achieves the best results.
Refer to caption(a)(b)(c)(d)(e)(f)
Figure 12: Comparison of visual results by different variants of excluding a certain loss in our method. (a) Input; (b) w.o. $L_{\textrm{R-SDF}}$; (c) w.o. $L_{\textrm{Norm}}$; (d) w.o. $L_{\textrm{Smo}}$; (e) w.o. $L_{\textrm{Shape}}$; f) Ours.
Refer to caption(a) Input(b) LPDC [50](c) Cadex [23](d) Ours00.15
Appendix 0.B More visual results
We present more visual results in this section and showcase the complete motion sequences in the Video Demo.
0.B.0.1 Comparisons with state-of-the-art methods.
We present additional visual results for comparisons with LPDC [50] and Cadex [23] on the AMA dataset [52] (Fig. 13), DT4D dataset [27] (Fig. 14) and DFAUST dataset [4] (Fig. 15). It is evident that our method outperforms other approaches.
Refer to caption(a) Input(b) Cadex [23](c) Ours00.0600.200.1
@@ Appendix 0.B More visual results
0.B.0.2 Performance on noisy data and partially missing data.
Furthermore, in Fig. 16, we showcase more visual results of our method applied to noisy data constructed from the DFAUST dataset [4] and partially missing data constructed from AMA dataset [52], respectively.
@@
Appendix 0.C Technical Details
In this section, we introduce the parameter settings and explain some definitions in the loss functions and evaluation metrics mentioned in the paper.
0.C.0.1 Parameter settings.
Reconstructing temporally consistent dynamic surfaces from point cloud sequences is challenging, especially for our unsupervised learning scenario. Therefore, we need multiple hyperparameters associated with the regularization terms for shrinking the large solution space of our unsupervised pipeline, we have uniformly set them to default values in all experiments and found that they generally work well for most examples, minimizing the need for additional parameter adjustments for new datasets. By default, in the calculation of the loss function (8), we set the number of the points for both $\widetilde{\mathbf{V}}_{k}$ in Eq. (9) and $\mathbf{Q}_{s}$ in Eq. (10) to be $10^{4}$. We set $\eta=0.1$ in Eq. (7), $\alpha=5.56$ in Eq. (9) and $\beta=50,\gamma=10^{2}$ in Eq. (10). The default weights setting in the loss function are listed in Tab. 5. In addition, we also showcase the number of iterations and training time in this table.
Table 5: Default parameter settings and runtime. For the learning template surface, we listed iterations and running time for the coarse/fine stage.
Learning template surface Temporal reconstruction
$5\times 10^{2}$ $10^{-3}$ $50$ $10^{3}$ / $5\times 10^{3}$ $10$ s/ $2$ mins $5\times 10^{2}$ $10^{-3}$ $10^{2}$ $10^{3}$ $1$ $10^{4}$ 30 mins
0.C.0.2 The details of loss functions.
The Chamfer Distance between $\mathbf{X}$ and $\mathbf{Y}$ is defined as
$\texttt{CD}_{\ell_{p}}(\mathbf{X},\mathbf{Y})=\frac{1}{|\widetilde{\mathbf{X}}|}\sum_{\mathbf{x}\in\widetilde{\mathbf{X}}}\left\|\mathbf{x}-\hat{\mathbf{y}}\right\|_{\ell_{p}}+\frac{1}{|\widetilde{\mathbf{Y}}|}\sum_{\mathbf{y}\in\widetilde{\mathbf{Y}}}\left\|\mathbf{y}-\hat{\mathbf{x}}\right\|_{\ell_{p}},$
@@ Appendix 0.C Technical Details
In Eqs. (3) and (8) of the paper, the normal consistency of $\mathbf{X}$ and ${\mathbf{Y}}$ is defined as
$\texttt{NC}_{\ell_{1}}(\mathbf{X},\mathbf{Y})=\frac{1}{|\mathbf{N}_{x}|}\sum_{\mathbf{n}_{x}\in\mathbf{N}_{x}}\left|1-|\langle\mathbf{n}_{x},\mathbf{n}_{\hat{y}}\rangle|\right|+\frac{1}{|\mathbf{N}_{y}|}\sum_{\mathbf{n}_{y}\in\mathbf{N}_{y}}\left|1-|\langle\mathbf{n}_{y},{\mathbf{n}}_{\hat{x}}\rangle|\right|,$
[…]
In Eq. (4), the SDF value $\texttt{SDF}_{{\rm IMLS}}(\mathbf{q},\mathbf{P}_{k^{*}})$ at $\mathbf{q}$ approximated through implicit moving least-squares is defined as
$\texttt{SDF}_{{\rm IMLS}}(\mathbf{q},\mathbf{P}_{k^{*}})=\frac{\sum_{\mathbf{p}_{j}\in\mathcal{N}(\mathbf{q})}\theta(\|\mathbf{q}-\mathbf{p}_{j}\|,{\zeta})\cdot\langle\mathbf{q}-\mathbf{p}_{j},\mathbf{n}_{j}\rangle}{\sum_{\mathbf{p}_{j}\in\mathcal{N}(\mathbf{q})}\theta(\|\mathbf{q}-\mathbf{p}_{j}\|,\zeta)},$
where $\mathcal{N}(\mathbf{q})$ denotes the set of nearest points in $\mathbf{P}_{k^{*}}$. $\theta(d,\zeta)=\exp(-d^{2}/\zeta^{2})$ is the weight of each point in $\mathcal{N}(\mathbf{q})$. Here we set $|\mathcal{N}(\mathbf{q})|=10$ and $\zeta=0.1$ by default.
0.C.0.3 Evaluation metrics.
@@
$\textrm{Recall}(\mathcal{M}_{k},\mathcal{M}_{k}^{\textrm{GT}},\epsilon)=\left|\left\{\mathbf{p}_{1}\in\widetilde{\mathbf{P}}_{k},\textrm{s.t.}\min_{\mathbf{p}_{2}\in\widetilde{\mathbf{P}}_{k}^{\textrm{GT}}}\|\mathbf{p}_{1}-\mathbf{p}_{2}\|<\epsilon\right\}\right|,$
$\textrm{Precision}(\mathcal{M}_{k},\mathcal{M}_{k}^{\textrm{GT}},\epsilon)=\left|\left\{\mathbf{p}_{2}\in\widetilde{\mathbf{P}}_{k}^{\textrm{GT}},\textrm{s.t.}\min_{\mathbf{p}_{1}\in\widetilde{\mathbf{P}}_{k}}\|\mathbf{p}_{1}-\mathbf{p}_{2}\|<\epsilon\right\}\right|.$
@@
We denote $\{\mathbf{P}^{\textrm{GT}}(k)\}_{k}$ as the temporally consistent corresponding points on the ground-truth surfaces. We set $|\mathbf{P}^{\textrm{GT}}(k)|=10^{5}$. In the first frame, we construct the index set of nearest points:
$\texttt{Index}=\{\rho_{i}|\mathbf{v}_{\rho_{i}}=\mathop{\arg\min}_{\mathbf{v}\in\mathbf{V}_{1}}\|\mathbf{v}-\mathbf{p}_{i}^{\textrm{GT}}(1)\|\},$
where $\mathbf{V}_{1}$ is the vertex set of $1$-st reconstructed mesh $\mathcal{M}_{1}$, $\mathbf{p}_{i}^{\textrm{GT}}(1)\in\mathbf{P}^{\textrm{GT}}(1)$. Following [36, 23], the Correspondences Error is defined as